OpenStudy (anonymous):

In calculus, how do you find the normal line equation when given the tangent line equation

OpenStudy (lgbasallote):

normal line means perpendicular line

OpenStudy (lgbasallote):

^does that help?

OpenStudy (anonymous):

normal line = \[-1/ f'(a) \]

OpenStudy (anonymous):

When you have a line in point-slope form,\[y-y_0=m(x-x_0)\]the unique line perpendicular to it is given as follows.\[y-y_0=-\frac{1}{m}(x-x_0)\]This is jus more of a line geometry problem than a calculus problem.

OpenStudy (anonymous):

@yakeyglee You actually have this in calculus.

OpenStudy (anonymous):

Thank you all :D

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

@beeqay - What I meant was that it wasn't a calculus technique.

OpenStudy (anonymous):

the normal line is the one that isn't abnormal...:)